I have done the following equations and I need to know "For what value is this equation undefined?"
6000/200-2x75%=120 million
6000/200-2x90%=300 million
6000/200-2x99%-3000million

Answer by stanbon(57224)


"For what value is this equation undefined?"
6000/200-2x75%=120 million
6000/200-2x90%=300 million
6000/200-2x99%-3000million
----
The denominator is (200-2x)
---
The fraction is undefined when 200-2x is zero.
---
Solve: 200-2x =0
---
2x = 200
x = 100
--------
Undefined when x = 100.
================================
Cheers,
Stan H.